OpenStudy (anonymous):

A group of people went to the movies. The price was $8 for Adults and $5 for Children. The total cost for 11 people was $67. How many adults and children went to the movies? (Write an equation(s) to solve the problem and show the steps to solve the problem.) please show how you solved it

OpenStudy (ranga):

Assume A adults and C children There were 11 people in all. A + C = ? $8 for adult ticket and $5 for children's ticket. What would be the total cost?

OpenStudy (anonymous):

would it make sense if i solve it like this? A=adults 8A+5*(11-A)=67 8A=money paid by adults.5*(11-A)=money paid by children. so equals 67 8A+55-5A=67 3A=12 x=4*(adults) Number of children =11-4 =7(children)

OpenStudy (ranga):

Yes. You will get the same answer. A + C = 11 ------ (1) 8A + 5C = 67 -------- (2) Multiply (1) by -5 -5A - 5C = -55 Add it to (2) 3A = 12 A = 4 Put A = 4 in (1) 4 + C = 11 C = 7